OpenStudy (anonymous):

so i have to find the equation of a line and i know how to do it but how do i know if my answer is right PPPPPPLLLEEEEAAAASSSSSEEEE HHHEEEELLLLLPPP tx

OpenStudy (anonymous):

Okay, so what's question? And, what did you get as the answer?

OpenStudy (anonymous):

@savygirl

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

p1=-2,3 p2=0,6

OpenStudy (anonymous):

and my answer was 3x-2y=-10

OpenStudy (anonymous):

and i dont know if its right if u could help i would greatly appreciate it

OpenStudy (anonymous):

you need the slope first it is \[\frac{6-3}{0-(-2)}=\frac{3}{2}\] then \[y-6=\frac{3}{2}x\] \[2y-12=3x\] \[3x-2y=-12\]

OpenStudy (anonymous):

everything right but your constant

OpenStudy (anonymous):

ok tx very much greatly appreciate it